

********************************************************************
UARS WEEKLY STATUS REPORT - 05 MAR 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE ended 6 days of daylight only operations on Friday (99064). In this
mode we did not take any stellar calibration observations but sucessfully
gathered our daily full solar data set. We are currently operating normally
taking both solar and stellar data during the orbit.
***********
SUSIM
***********
SUSIM was turned off on Feb. 26 and will resume operations on March 5.
Reprocessing of the SUSIM data to data version 20 continues.
***********
PEM
***********
PEM is active under PMON control. Energy processor #6 for the
HEPS-1 unit has been found to not activate correctly. Data affected
are from the upper energy EE channel for Head-2 of HEPS-1 (that is the
+15 deg electrons above 300 keV). Tests are being conducted on this
energy processor to assure that it is operating correctly. Measurements
from this sensor will be activated as soon as possible. The PEM team
and FOT are convinced that the entanglement of energy processor #6
is related to OBC control and not a failure of that energy processor.
Early attempts at adjusting OBC command timing have not been successful.
The FOT is attempting to refine their diagnostics and confine the
problem.
While under PMON control, the PEM instrument computes energy deposition
from the HEPS unit.
Between Feb 22 and Feb 26, the PEM X-ray unit 1 was active. During this
time only eight of 16 AXIS pixels were active and provided atmospheric energy
deposition over half of the satellite track. The AXIS unit emphasized
energy deposition from electrons (inverted from the measured x-ray spectrum)
measured from the cold side of the spacecraft, reaching the highest sampled
latitudes over the north pole.
***********
HALOE
***********
HALOE continues successful operations with all engineering parameters
within their nominal operating range.
***********
MLS
***********
MLS was reactivated successfully on Feb 27 and operated though Mar 4.
***********
HRDI
***********
HRDI is currently operating during daytime only, and is functioning normally.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF notified Instrument PI software representatives regarding updated
LEVEL0 Y2K files. This update also includes Attitude EXTRSC and SSPPGIMBALS
files that were generated against the LEVEL0 Y2K files. Please contact PAC
if you need the location.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
7 days ATT (7.3) RSS (Y2K) (7R-7C=0L)
3 days CLA (7.0) ENG (L1-L3) (3R-3C=0L)
3 days HAL (4.81) PROD (TEST L1-L3) (3R-3C=0L)
50 days PEM (3.32) PROD (OA-L2) (760R-180C=580L)
1 day SEN (5.0) PROD (1R-1C=0L)
189 days SUS (7.0) PROD (668R-256C=412L)
"Solstice Wall Clock time increase.
Attempts to determine the cause of the "insufficient virtual memory"
error which occurs when running the debug version of SOLSTICE on the
Alpha have been unsuccessful. Plan to run the SOLSTICE job, in debug,
on the VAX (which is not how it is run natively but the same database
is used). Currently setting up for VAX run."
CDHF pending software upgrades: DEC AXP Fortran 6.5, DEC AXP OPEN
VMS 6.2, OpenIngres.
***********
FOT
***********
This report covers 20 February 1999 (Orbit 40676, GMT Day 051 through 26
February 1999.
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 053/0157 40707 AXIS1 On
HALOE 054/2253 40735 On
SUSIM 057/2301 40780 Off
ACRIM 057/2301 40780 Off
PEM 057/2301 40780 AXIS1 Off
The beta angle ranged between the angles of 6.1 degrees to 22.9 degrees for
this report period. The beta angle is now increasing toward a maximum
angle of +53 degrees on 10 March 1999 (DOY 069).
A Reverse-to-Forward Yaw Around maneuver (number 76) was successfully
executed in Orbit 40686 beginning 051/15:27:48 and completing at
051/16:04:58 on 20 February 1999. The SSPP Instruments (ACRIM, SOLSTICE,
SUSIM) were enabled as well as HRDI and PEM in day only operations. Due to
the SA position at 269 degrees, the effective night length of the
spacecraft after the maneuver was approximately 56 minutes. Post maneuver
analysis of data indicate that during the Yaw Around, the end-of-night load
bus voltage was 24.8 volts and the average depths of discharge (DODs) on
Batteries 2 and 3 were approximately 22.0 and 22.0 percent respectively.
The next (forward-to-reverse) YAW is scheduled for 28 March 1998 (DOY 087).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.9 to 1.3 degrees C. Additional battery performance data
for each battery on 19 February 1999 and 26 February 1999 are provided in
Appendix D.
The clock error ranged from -5.2 to 5.4 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 03
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 00
% ACRIM 01
% HALOE 02
% PEM 04
% SUSIM 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 00
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 99/019 - 2287.496951
Transponder B: 99/020 - 2287.499045
GROUND SYSTEM OPERATION
No (0) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 99/056 show the observatory data loss to be
22 hours, 46 minutes, 27.604 seconds (an increase of 01.632 seconds since
last report period). This is a 0.0350 percent data loss which equals a
99.9650 percent data capture for the mission.
The increase of 01.632 seconds occurred over three (3) days:
7 0.224 seconds on Day 050 due to data hit,
7 0.800 seconds on Day 054 due to data hit, and
7 0.608 seconds on Day 056 due to data hit.
The WINDII Principle Investigator, has requested UARS to operate with the
cold side rolled down 0.5 degrees for the period 05 March 1999 through 16
March 1999 (DOY 064 through 075). Planning is underway.
FOT requested an updated set of weekly planning aids from FDF on 25
February to support the request from WINDII Operations to perform a 0.5
degree roll offset maneuver on 04 March (lasting through 16 March),. If
the FOT waited until the next normal delivery of weekly planning aids (on
03 March), the instrument operations teams would not have enough time to
plan for the maneuver. Therefore, the additional set of planning aids
(which include the roll effects) was necessary.
No additional Y2K "test" 8-week planning aids were received by the FOT this
week. The FOT is still awaiting the delivery of two "corrected" planning
aids to complete the test set. These two remaining files will most likely
be delivered to the FOT next week, after which time the FOT will process
the files and make them available to the MPG and Instrument Operations
Teams.
The UARS Test and Training Simulator (UTTS) remains non-operational. The
estimated time to repair the UTTS and return it to operation is unknown.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continues to support the PEM team with diagnosis of the Energy
Processor 6 status error. PEM EP6 is still not coming up nominally under
PMON control. Further timing tests suggest that EP6 is functioning after
normal activation sequences, and during a "time-expanded" ground commanded
version of the PMON activation sequence. This suggests a timing conflict
in the PMON day-only activation sequence. The FOT will conduct additional
tests to identify where in the turn-on sequence the conflict occurs in the
coming week.
The FOT continues to investigate the feasibility of a new PMON action to
change battery V/T curves based on loading to mitigate battery risks in sun
pointing safehold.
YEAR 2000 OPERABILITY SUPPORT - No change
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
Planning continues for the simulator testing.
UARS RE-ENGINEERING - TELEMETRY PROCESSING - No change
MEETINGS
FOT participated in a meeting on 23 February 1999 concerning the
development of an Operational Concept Document for the proposed Integrated
Mission Operation Complex (IMOC).
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver (forward-to-reverse) is scheduled for 28
February 1999 (DOY 087).
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 26 February 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
40707,8 053/0157 99-072 PEM AXIS1 ON
40732 054/1751 99-073 PEM Amend DAY ONLY Procedure
40733 054/1934 99-074 PEM Amend DAY ONLY Procedure
40735 054/2253 99-075 HALOE Power ON
40741 055/0812 99-076 HALOE Enabled Mission Unique Bit
40780 057/2301 99-077 SUSIM Power OFF
40780 057/2301 99-078 ACRIM Power OFF
40780 057/2301 99-079 PEM Power OFF, AXIS
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 26 February 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
PEAXTHERM6E Y-HI AXIS2 AMP BOX 40711-780
HASMIRRT Y-HI HALOE SEC. MIRROR TEMP 40738
PEMBIAS02E R-HI Sensor 2 Grid Bias Monitor 40771
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 26 February 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
-No clock adjusts this report period
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 26 February 1999
19 Feb 1999 - GMT Day 050 Beta = 6.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 0.0/-22.4 -3.68/-4.94 +0.0/ +0.0 16.8 0.000 0.0
2 +67.2/ 0.0 +4.83/+3.87 +31.2/-14.4 25.6 1.016 20.8
3 +89.6/-5.6 +3.87/+2.60 +30.4/-14.4 25.4 1.017 20.4
26 Feb 1999 - GMT Day 057 Beta = 22.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -5.6/-33.6 -3.05/-4.31 +0.0/ +0.0 16.8 0.000 0.0
2 +61.6/ 0.0 +5.47/+4.51 +31.6/-14.0 25.4 1.024 21.2
3 +67.2/-16.8 +4.51/+3.23 +31.2/-15.2 25.4 1.024 21.0
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 26 February 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-007 FOT 40577 044/1019 N/A no lock on 32k OBC dump
99-008 FOT 40589 045/0419 20585 "No UARS RF, no TDE trk"
99-009 FOT 40590 045/0650 14170 Generic Late Acquisition
99-010 FOT 40636 048/0750 N/A GCMR to 32k on Q invalid
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-006 FOT see list Feb 99 14170 Generic Late Acq
SEVNTFW99/09